Question 32
ProbabilityA bag contains 8 red, 4 blue and 2 green identical balls. Two balls are drawn randomly from the bag without replacement. Find the probability that the balls drawn are red and blue.
- A.12/91
- B.16/91
- C.30/91
- D.32/91
Show the answer and explanation
Answer: D (32/91)
Explanation
Red then blue or blue then red: (8/14)(4/13) + (4/14)(8/13) = 32/182 + 32/182 = 32/91.